Just minus For minus YouJust−For−You Hardware Consultants purchased a building for $ 540 comma 000$540,000 and depreciated it on a​ straight-line basis over a 3030​-year period. The estimated residual value is $ 120 comma 000$120,000. After using the building for 15​ years, Just minus For minus YouJust−For−You realized that wear and tear on the building would wear it out before3030years and that the estimated residual value should be$ 106 comma 000$106,000.Starting with the 16th​ year,Just minus For minus YouJust−For−Youbegan depreciating the building over a revised total life of2020years using the new residual value. Journalize depreciation expense on the building for years 15 and 16. ​(Record debits​ first, then credits. Select the explanation on the last line of the journal entry​ table.)

Begin by journalizing the depreciation on the building for year 15.
Answer #1
Journal entry
Particulars Debit Credit
Year 15 depreciation expenses on building 14000
accumulated depreciation on building 14000
depreciation expenses = purchase cost-residual value/estimated life
depreciation expenses = 540000-120000/30 = 14000
Book value at the end of year 15
540000-(14000*15) = 330000
Remaining life = 20-15 = 5
depreciation = (330000-106000)/5 = 44800
Journal entry
Particulars Debit Credit
Year 16 depreciation expenses on building 44800
accumulated depreciation on building 44800

Please give a thumbs up if it is helpful & let me know if any doubt